The electric field above a uniformly charged nonconducting sheet is E .If the non conducting sheet is now replaced by a conducting sheet ,with charge same as before ,new electric field at same point is=

Grade:12th pass

1 Answers

Susmita
5 years ago
For nonconduction sheet electric field
E=\frac{\sigma}{2\epsilon_{0}}
For conducting sheet electric field
E_{c}=\frac{\sigma}{\epsilon_{0}}
So Ec=E/2
.............................................................